x - 10 = -6x + 10
Solution: x = 12
Please post your question completely .
Is it solve for x or check the solution x = 12
I am answering it for the solve for x .

Answer by grishma.kshatriya(27)

Solution :
x - 10 = -6x + 10
We will combine the like terms ,
Add 6x on both the side we get ,
x + 6x - 10 = 10
7x - 10 = 10
Add 10 on both the sides ,
7x = 0
x = 0
Thank you .